OpenStudy (anonymous):

The numbers twenty, twenty one, twenty nine. form a Pythagorean triple. Which other set of numbers is a Pythagorean triple? twenty one, twenty two, thirty forty, forty two, fifty eight sixty, sixty six, eight seven seventy, eighty four, one hundred sixteen

OpenStudy (anonymous):

B

geerky42 (geerky42):

Give help, not answers... @SoulSand http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

3, 4, 5 21/3= 22/4= 30/5= are they all equal? If yes then that is your answer. 40/3= 42/4= 48/5= are they all equal? If yes then that is your answer. 60/3= 66/4= 87/5= are they all equal? If yes then that is your answer. 70/3= 84/4= 116/5= are they all equal? If yes then that is your answer.